OpenStudy (anonymous):

In a class of 42 students, 18 are boys. Two girls get transferred to another school. What percent of students remaining are girls?

OpenStudy (anonymous):

How many girls were there in the class originally?

OpenStudy (anonymous):

42-18=24 so we have 24 girls so 24 2 go away co we have 22 22/42=?

OpenStudy (anonymous):

@DejaVu5

OpenStudy (anonymous):

When two students leave, there are no longer 42 students in the class.

OpenStudy (anonymous):

oh right haha my bad 22/40....... nice catch @ospreytriple

OpenStudy (anonymous):

weeeeelllllll.... divide it then multiply by 100 you'll get your percentage.......

OpenStudy (anonymous):

do you mind drawing it out for me?

OpenStudy (anonymous):

there's really nothing to draw, i mean its all there.... want me to explain the steps?

OpenStudy (anonymous):

okay so first off we want to find out how many girls are in the class, so we subtract the boys (18) from the total (42), thus we get the number of girls (24). so so if two girls transferred (-2) we subtract 2 from both the total (42) and the number of girls (24). we get 22 girls and 40 students in the room, so to get the percentage we divide 22 by 40, where we get 0.55. then we multiply 0.55 but 100 to get the percentage: 0.55*100= 55 so the percentage of girls in the class is 55%

OpenStudy (anonymous):

@DejaVu5